OpenStudy (anonymous):

Evaluate cot 45° without using a calculator by using ratios in a reference triangle.

OpenStudy (madhu.mukherjee.946):



OpenStudy (anonymous):

\[\theta = 45 \] right?

OpenStudy (madhu.mukherjee.946):

yeah

OpenStudy (madhu.mukherjee.946):

so now tan theta=perpendicular/base

OpenStudy (anonymous):

okay so i need to get cot 45°, what two would i have to use

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 1 }\]